# Neural Posterior Estimation for Cataloging Astronomical Images with Spatially Varying Backgrounds and Point Spread Functions

### v1

@@ Abstract
Neural posterior estimation (NPE), a type of amortized variational inference, is a computationally efficient means of constructing probabilistic catalogs of light sources from astronomical images. To date, NPE has not been used to perform inference in models with spatially varying covariates. However, ground-based astronomical images have spatially varying sky backgrounds and point spread functions (PSFs), and accounting for this variation is essential for constructing accurate catalogs of imaged light sources. In this work, we introduce a method of performing NPE with spatially varying backgrounds and PSFs. In this method, we generate synthetic catalogs and semi-synthetic images for these catalogs using randomly sampled PSF and background estimates from existing surveys. Using this data, we train a neural network, which takes an astronomical image and representations of its background and PSF as input, to output a probabilistic catalog. Our experiments with Sloan Digital Sky Survey data demonstrate the effectiveness of NPE in the presence of spatially varying backgrounds and PSFs for light source detection, star/galaxy separation, and flux measurement.
@@ 1 Introduction
NPE has numerous advantages for cataloging. First, it is a likelihood-free method in that it does not require evaluation of the conditional likelihood function; instead, it suffices to draw samples from the joint distribution. This means even complex simulators, such as GalSim (Rowe et al., 2015) and PhoSim (Peterson et al., 2015), can be used to encode the conditional likelihood. Second, NPE enjoys important theoretical guarantees that are not shared by traditional (ELBO-based) approaches to amortized variational inference, which minimize the reverse Kullback-Leibler (KL) divergence, that is, the KL divergence from the posterior approximation to the posterior. NPE instead minimizes the forward Kullback-Leibler divergence and marginalizes over nuisance latent variables implicitly by simply excluding them from the objective function (Ambrogioni et al., 2019). Further, with reasonable conditions, NPE is globally convergent, whereas other types of variational inference often converge to shallow local optima (McNamara et al., 2024).
To date, NPE has not been used to catalog astronomical images with spatially varying covariates, such as backgrounds and PSFs, despite the importance of doing so. In this work, we propose a statistical model that treats the background and PSF of images, along with the latent properties of imaged light sources, as random variables in a Bayesian model (Section 4). In this model, the backgrounds and PSFs follow the empirical distribution of backgrounds and PSFs estimated by an existing survey.
Next, we propose an NPE procedure to train an inference network to estimate the parameters used to generate the data, such as the number of objects in the image and their positions and fluxes (Section 5). One challenge lies in designing an inference network capable of extracting features on two vastly different spatial scales. Light sources usually cover hundreds of pixels (e.g., a 20$\times$20-pixel patch), while point spread functions (PSFs) and backgrounds vary slowly and exhibit spatial dependencies that span millions of pixels and multiple images. To accurately infer the properties of a light source, an inference network must attend to the broader context, to learn about the background and PSF, and to the highly localized image region that pertains to the particular light sources. A sophisticated network design would be needed to operate on such different scales. If multiple images are required to infer the broader context, a sophisticated data loading scheme would be needed as well.
@@
Once trained, this inference network takes images, survey-estimated backgrounds, and survey-estimated PSFs as input. It outputs a posterior approximation in a single forward pass, without any additional training. We demonstrate the robustness of our method by applying it to images from the Sloan Digital Sky Survey (SDSS) and show that amortized inference can be a useful tool even when the background and PSF vary across images (Sections 6 and 7). Finally, we discuss the limitations of our approach and propose possible extensions to improve the method (Section 8).
@@ 2 The SDSS Image-Processing Pipeline
The semi-synthetic data that we generate for training and validating our method is based on data from the Sloan Digital Sky Survey (SDSS). The SDSS is a spectroscopic and astronomical imaging survey designed to study the redshift of extragalactic content. SDSS data have been used for a variety of research objectives, including investigations of galaxy formation and evolution, quasars, and the evolution of dark energy. Although our method can be extended for use with data from any survey, the SDSS is a suitable dataset for validating the proposed method because of its well-understood survey characterization and longevity. We used both SDSS images and the SDSS estimates of background intensity and PSF for these images.
@@
PSF estimation. The SDSS pipeline, known as photo, models PSFs using the Karhunen-Loéve decomposition to decompose stars brighter than 20 magnitude into eigenimages and retains the first three components (Lupton et al., 2005). The pipeline then fits a model of the radial profile of the PSF composed of two Gaussians and a power-law model. At radius $r$ from the center, the PSF is given by
$\mathbf{\Pi}(r)$ $=\frac{\exp{\left(-\frac{r^{2}}{2\sigma_{1}^{2}}\right)}+b\exp{\left(-\frac{r^{2}}{2\sigma_{2}^{2}}\right)}+p_{0}\left(1+\frac{r^{2}}{\beta\sigma_{p}^{2}}\right)^{-\frac{\beta}{2}}}{1+b+p_{0}},$
where $\sigma_{1}$ and $\sigma_{2}$ are standard deviations for the two Gaussians, $b$ is the ratio of the second Gaussian to the first at the origin, $\sigma_{p}$ is the width parameter for the power-law model, $p_{0}$ is the value of the power law at the origin, and $\beta$ is the slope of the power law. Examples of PSFs from two SDSS fields are shown in Figure 1. There is a visible difference in the PSF sizes for the two fields. The background and PSF variation for multiple nightly runs of the Sloan Digital Sky Survey (SDSS) are illustrated in Figure 2. The wide variation in both the PSF size and background level in the SDSS dataset further illustrates its suitability as a dataset for validation of our method.
@@ 3 Methodology
Figure 3: Overview of the inference procedure. The catalog (light grey, left) follows a parametric prior, whereas the background and PSF follow SDSS empirical distributions; they are sampled by selecting an SDSS field uniformly at random. Together, a catalog, PSF, and background define a sky image that is generated by GalSim, which adds Poisson shot noise. The inputs to the inference network are the image, the background, and the PSF parameter encoding. The inference network outputs a posterior approximation of the catalog conditioned on the image, background, and PSF (light gray, right).
@@ 4.1 Prior Distribution
where $H$ and $W$ are the height and width of the image in pixels and $\mu=6.25\times 10^{-4}$, to resemble the typical per-pixel source density in SDSS. For each $s=1,\ldots,|z|$, the properties of light source $s$, denoted $z_{s}$, include position, source type (star or galaxy), flux, and galaxy shape. They are sampled according to the specification in Table 1, which is fitted to earlier catalogs of SDSS.
We set $p(\gamma,\Pi)$ to the empirical joint distribution of backgrounds and PSFs in a diverse subset of SDSS fields. For simplicity, for each SDSS field, we consider only the modeled PSF for the field center, rather than selecting the PSF for a random position within each field. Sampling a background and PSF entails choosing an SDSS field uniformly at random and then returning the background and PSF from the photo pipeline for that field.
@@ 4.2 Conditional Likelihood
Astronomical images record the number of photons detected in a given pixel. The intensity of each pixel depends on three factors: the properties of the light sources in the image (e.g., positions, source types, fluxes, and shapes), the background intensity, and the point spread function.
Given a catalog $z$, background $\gamma$, and PSF $\Pi$, we model the photon count $x_{ij}$ at pixel $(i,j)$ as a Poisson process with a rate $\lambda_{ij}$ that depends on $z$, $\gamma$, and $\Pi$:
@@
where $\gamma_{ij}$ is the background intensity for pixel $(i,j)$ and $\xi_{ij}$ is a deterministic function that maps a source’s properties to its expected contribution to pixel $(i,j)$ based on the PSF. This deterministic function is encoded by the GalSim image simulator (Rowe et al., 2015). Since the photon counts $x_{ij}$ at pixel $(i,j)$ are expected to be high, we use a Gaussian approximation to the Poisson distribution:
$x_{ij}\sim\mathcal{N}(\lambda_{ij},\lambda_{ij})$ (4)
Finally, we subtract the background from the image to standardize the mean intensity across images:
$x_{ij}:=x_{ij}-\gamma_{ij}$
@@ 5 Neural Posterior Estimation: Bayesian Inference with a Neural Network
Given an astronomical image $x$, background $\gamma$, and PSF $\Pi$, we aim to infer the posterior distribution of the catalog, $p(z\mid x,\gamma,\Pi)$. This posterior is intractable as it involves integrating over all possible catalogs. Therefore, instead of inferring the posterior exactly, we approximate the posterior using neural posterior estimation (NPE). NPE aims to select a variational approximation $q_{\phi}$ indexed by parameters $\phi\in\Phi$ that minimizes the forward Kullback-Leibler (KL) divergence:
$q_{\phi^{*}}=\operatorname*{arg\,min}_{\phi\in\Phi}D_{\rm KL}(p\,\|\,q_{\phi}).$
We accomplish this by training a neural network to map $(x,\gamma,\Pi)$ to a parameterized representation of $q(z\mid x,\gamma,\Pi)$, which we describe further in Section 5.1.
Our network is trained using the simulated data described in Section 4. The weights of the neural network are learned through stochastic optimization to minimize the negative log likelihood:
$\mathcal{L}:=-\log q(z\mid x,\gamma,\Pi)$
After the network has been trained, evaluating the posterior for a given data point $(x,\gamma,\Pi)$ just involves a single forward pass through the neural network.
@@ 5.1 The Variational Family
The probabilistic structure of our variational distribution, which factorizes over small spatial regions called “tiles,” matches that of Liu et al. (2023b). In particular, let $H^{\prime}=\lceil H/4\rceil$ and $W^{\prime}=\lceil W/4\rceil$. For $h=1,\ldots,H^{\prime}$ and $w=1,\ldots,W^{\prime}$, let image tile $T_{h,w}:=[h,h+4)\times[w,w+4)$ refer to particular 4$\times$4-pixel region. Collectively, the image tiles are a partition of the region imaged by $x$, expressed in pixel coordinates. For a catalog $z$, let $z_{h,w}\subset z$ denote the cataloged light sources whose centroids are contained in $T_{h,w}$.
To approximate $p(z\mid x,\gamma,\Pi)$, we use a family of variational distributions that factorize over tiles. In particular, we set
$q(z\mid x,\gamma,\Pi):=\prod_{h=1}^{H^{\prime}}\prod_{w=1}^{W^{\prime}}q(z_{h,w}\mid x,\gamma,\Pi),$
where the probabilistic structure of $q(z_{h,w}\mid x,\gamma,\Pi)$, which we refer to as the per-tile variational distribution, is given in Table 2. For each tile, this distribution is governed by eight distribution parameters. For a tile with index $(h,w)$, these distributional parameters are outputted in spatial position $(h,w)$ by a fully convolutional inference network that takes $x$ as input. We describe its architecture next.
@@ 5.2 Neural Network Architecture
The input to our inference neural network is an $80\times 80$-pixel image (i.e., $H=80$ and $W=80$) of stars and galaxies with a given background and PSF. To approximate the posterior for the catalog, we also input the image background and our PSF model parameter encoding (see Figure 3).
Providing the PSF encoding to the network is more complex than providing the background. One option is to explicitly account for the PSF in a convolutional neural network (CNN) by appending an image of the PSF as another channel, as done in Lanusse et al. (2021), when estimating galaxy morphologies from postage stamp images. But, this may not generalize well to images with a larger footprint, where we would ideally want the input accompanying each pixel to encode the entirety of the PSF.
One potential remedy is to apply a deconvolution to the input image using the known PSF, appending the deconvolved image to the input. However, deconvolution is an ill-posed inverse problem. We tested this potential remedy and found that it scarcely improves the performance of an inference network without PSF information.
We propose an alternative approach that makes use of a low-dimensional encoding of the PSF as follows. For a $d$-dimensional representation, we concatenate $d$ separate input feature maps to the image, where each input feature map consists of the same value at each position and has the same height and width as the image. This guarantees that the relevant PSF information can be accessed in the same way by a CNN at each pixel position. For simplicity, we use the six PSF parameters from the SDSS pipeline (described in Section 2), as our low-dimensional encoding. We discuss the possibility of using an auxiliary autoencoder as a more sophisticated method of obtaining this encoding in Section 8. The inputs to the network are either (i) a single-channel image, (ii) an image and background, or (iii) an image, background, and six additional channels, one for each of the PSF parameters.
6 Design of Numerical Experiments
Through numerical experiments, we aim to assess the impact of incorporating background and PSF information in our networks that predict the centroid, flux, and type of each light source in an image. Specifically, we wish to quantify the robustness of our method to model misspecification, where the inference network has been trained with data sampled from the model that may differ slightly from the distribution that produces a given observation. The code for our experiments is publicly available at https://github.com/prob-ml/bliss/.

### v2

@@ Abstract
Neural posterior estimation (NPE), a type of amortized variational inference, is a computationally efficient means of constructing probabilistic catalogs of light sources from astronomical images. To date, NPE has not been used to perform inference in models with spatially varying covariates. However, ground-based astronomical images exhibit spatially varying sky backgrounds and point spread functions (PSFs), and accounting for this variation is essential for constructing accurate catalogs of imaged light sources. In this work, we introduce a novel NPE-based cataloging method that trains an inference network with semi-synthetic astronomical images generated using PSF and backgrounds sampled from from the Sloan Digital Sky Survey. In experiments with semi-synthetic images, we evaluate the method on key cataloging tasks: light source detection, star/galaxy separation, and flux measurement. A “generalist” inference network—trained with diverse PSFs and backgrounds—performs as well as a “specialist” network even when both are evaluated on the specialist’s particular PSF/background combination. This result suggests that a single NPE network can generalize across spatial variations, eliminating the need for retraining on each observational condition.
@@ 1 Introduction
NPE has numerous advantages for cataloging. First, it is a likelihood-free method in that it does not require evaluation of the conditional likelihood function; instead, it suffices to draw samples from the joint distribution. This means even complex simulators, such as GalSim (Rowe et al., 2015) and PhoSim (Peterson et al., 2015), can be used to encode the conditional likelihood. Second, NPE enjoys important theoretical guarantees that are not shared by traditional (ELBO-based) approaches to amortized variational inference, which minimize the reverse Kullback-Leibler (KL) divergence, that is, the KL divergence from the posterior approximation to the posterior. NPE instead minimizes the forward Kullback-Leibler divergence and marginalizes over nuisance latent variables111 Nuisance latent variables are unobserved model parameters that are necessary for accurate forward modeling but are not themselves of direct scientific interest. In our work, the PSF and background intensity levels are nuisance variables: they reflect properties of the instrumentation and observing conditions rather than of the astronomical objects we aim to catalog. implicitly by simply excluding them from the objective function (Ambrogioni et al., 2019). Further, with reasonable conditions, NPE is globally convergent, whereas other types of variational inference often converge to shallow local optima (McNamara et al., 2024).
To date, NPE has not been used to catalog astronomical images with spatially varying covariates, such as backgrounds and PSFs, despite the importance of doing so. In this work, we propose a statistical model that treats the background and PSF of images, along with the latent properties of imaged light sources, as random variables in a Bayesian model (Section 4). In this model, the backgrounds and PSFs follow the empirical distribution of backgrounds and PSFs estimated by an existing survey. In principle, background samples could come from any source, including both the prior and posterior of sophisticated structured models of Saydjari & Finkbeiner (2022) and Feder et al. (2023). Here, however, we focus on the empirical distribution of background estimates produced by the Sloan Digital Sky Survey’s photometric pipeline, which exhibits more gradual spatial variation than samples from these structured models.
Additionally, we propose an NPE procedure to train an inference network to estimate the parameters used to generate the data, such as the number of objects in the image and their positions and fluxes (Section 5). One challenge lies in designing an inference network capable of extracting features on two vastly different spatial scales. Light sources usually cover hundreds of pixels (e.g., a $\times20$-pixel patch), while point spread functions (PSFs) and backgrounds vary slowly and exhibit spatial dependencies that span millions of pixels and multiple images. To accurately infer the properties of a light source, an inference network must attend to the broader context, to learn about the background and PSF, and to the highly localized image region that pertains to the particular light sources. A sophisticated network design would be needed to operate on such different scales. If multiple images are required to infer the broader context, a sophisticated data loading scheme would be needed as well.
@@
Once trained, this inference network takes images, survey-estimated backgrounds, and survey-estimated PSFs as input. It outputs a posterior approximation in a single forward pass, without any additional training. We demonstrate the effectiveness of our method by applying it to semi-synthetic images designed to mimic images from the Sloan Digital Sky Survey (SDSS) Section 6). Finally, we discuss the limitations of our work and propose extensions to improve the proposed method (Section 7).
@@ 2 The SDSS Image-Processing Pipeline
The semi-synthetic data that we generate for training and validating our method is based on data from the Sloan Digital Sky Survey (SDSS). The SDSS is a spectroscopic and astronomical imaging survey designed to study the redshift of extragalactic content. SDSS data have been used for a variety of research objectives, including investigations of galaxy formation and evolution, quasars, and the evolution of dark energy. While our method is adaptable to data from any survey, the SDSS offers an ideal test case due to its well-characterized properties and long operational history.
In SDSS, photometric completeness for galaxies decreases notably around 22 r-band magnitude. In DR7, S11 field, comparisons with the deeper COMBO‑17 survey indicate that SDSS recovers only about 50% of galaxies at 22 r-band magnitude (SDSS Collaboration, 2006). Further, Johnston et al. (2007) introduced formal magnitude‑completeness tests, which statistically validate that SDSS reaches its claimed faint limits. Although SDSS attains a $5\sigma$ point-source sensitivity of 22.7 r-band magnitude, this pertains solely to isolated stars; extended galaxy completeness declines significantly by 22 magnitude, marking the onset of a regime characterized by lower signal-to-noise and reduced detection efficiency.
In SDSS, star/galaxy separation performance has been benchmarked in multiple ways. The original Photo pipeline classified sources based on the difference between PSF and model magnitudes, aiming to distinguish point-like and extended sources (Lupton et al., 2001). This morphological separation is reliable down to $r\sim 21$, beyond which classification accuracy diminishes due to increasing similarity between compact galaxies and stars (Scranton et al., 2002; Yahata et al., 2007). Ball et al. (2006) introduced a machine learning approach using decision trees trained on a spectroscopic sample from SDSS DR3. Their classifier achieved galaxy completeness and efficiency of 98.9% and 98.3%, respectively, and star completeness and efficiency of 91.6% and 93.5%, respectively, within the magnitude range of $r<17.77$. They reported consistent performance to $r\sim 20$, with decreasing reliability at fainter magnitudes.
In addition to these benchmarks, the SDSS pipeline’s estimates of background intensity and PSF are important for our work.
@@
PSF estimation. The SDSS pipeline, known as photo, models PSFs using the Karhunen-Loéve decomposition to decompose stars brighter than 20 magnitude into eigenimages that retain the first three components (Lupton et al., 2005). The spatial variation of a PSF across an image is captured by modeling the coefficients of these components as low-order polynomials in position. From this spatially varying, pixelized PSF model, the pipeline then fits a smooth, radially symmetric mixture model composed of two Gaussians and a power-law tail, with six parameters in total: $\sigma_{1}$ and $\sigma_{2}$ are standard deviations for the two Gaussians, $b$ is the ratio of the second Gaussian to the first at the origin, $\sigma_{p}$ is the width parameter for the power-law model, $p_{0}$ is the value of the power law at the origin, and $\beta$ is the slope of the power law. At radius $r$, this PSF model is
$\mathbf{\Pi}(r)$ $=\frac{\exp{\left(-\frac{r^{2}}{2\sigma_{1}^{2}}\right)}+b\exp{\left(-\frac{r^{2}}{2\sigma_{2}^{2}}\right)}+p_{0}\left(1+\frac{r^{2}}{\beta\sigma_{p}^{2}}\right)^{-\frac{\beta}{2}}}{1+b+p_{0}}.$
Figure 1 shows PSFs for two SDSS fields. There is a visible difference in their sizes. The background and PSF variation for multiple nightly runs of the Sloan Digital Sky Survey (SDSS) are illustrated in Figure 2. The wide variation in both the PSF size and background level in the SDSS dataset further illustrates its suitability as a dataset for validation of our method.
@@ 3 Methodology
Figure 3: Overview of the inference procedure. The catalog (light grey, left) follows a parametric prior, whereas the background and PSF follow SDSS empirical distributions; they are sampled by selecting an SDSS field uniformly at random. Together, a catalog, PSF, and background define a sky image that is generated by GalSim, which adds Poisson shot noise. The inputs to the inference network are the image and the PSF. The inference network outputs a posterior approximation of the catalog conditioned on the image and PSF (light gray, right).
@@ 4.1 Prior Distribution
where $H$ and $W$ are the height and width of the image in pixels and $\mu=6.25\times 10^{-4}$, to resemble the typical per-pixel source density in SDSS. For simplicity, $\mu$ is fixed in our work; modeling $\mu$ as random could better represent spatial variation in source density. For each $s=1,\ldots,|z|$, the properties of light source $s$, denoted $z_{s}$, include position, source type (star or galaxy), flux, and galaxy shape. They are sampled according to the specification in Table 1, which is fitted to earlier catalogs of SDSS.
We set $p(\gamma,\Pi)$ to the empirical joint distribution of backgrounds and PSFs in a diverse subset of SDSS fields. Sampling a background and PSF entails choosing an SDSS field uniformly at random and then returning the background and PSF from the photo pipeline for that field. An alternative choice would be to set $p(\gamma,\Pi):=p(\gamma)p(\Pi)$, that is, the product of marginals, and to sample backgrounds and PSFs independently from SDSS. This alternative prior ignores potential correlation between background and PSF, and thus produce a overdispersed prior. Overdispersed priors are widely used in statistics, and ignoring correlation in the prior does prevent recovering correlation in the posterior. However, because we have a ample supply of paired backgrounds and PSFs, we do not explore this-data efficient alternative.
For simplicity, for each SDSS field, we consider only the modeled PSF for the field center, rather than selecting the PSF for a random position within each field.
@@ 4.2 Conditional Likelihood
Astronomical images record the the intensity of electromagnetic radiation in a given pixel. The intensity of each pixel depends on three factors: the properties of the light sources in the image (e.g., positions, source types, fluxes, and shapes), the background intensity, and the point spread function.
Given a catalog $z$, background $\gamma$, and PSF $\Pi$, we model the count of photoelectrons $x_{ij}$ at pixel $(i,j)$ as a Poisson process with a rate $\lambda_{ij}$ that depends on $z$, $\gamma$, and $\Pi$:
@@
where $\gamma_{ij}$ is the background intensity for pixel $(i,j)$ and $\xi_{ij}$ is a deterministic function that maps a source’s properties to its expected contribution to pixel $(i,j)$ based on the PSF. This deterministic function is encoded by the GalSim image simulator (Rowe et al., 2015). Since the background intensity $\gamma_{ij}$ is always at least several hundred, we use a Gaussian approximation to the Poisson distribution:
$x_{ij}^{\prime}\sim\mathcal{N}(\lambda_{ij},\lambda_{ij})$ (4)
Finally, we subtract the background from the image:
$x_{ij}:=x_{ij}^{\prime}-\gamma_{ij}$
@@ 5 Neural Posterior Estimation: Bayesian Inference with a Neural Network
Given a sky-subtracted astronomical image $x$ and PSF $\Pi$, we aim to infer the posterior distribution of the catalog, $p(z\mid x,\Pi)$. This posterior is intractable as it involves integrating over all possible catalogs. Therefore, instead of inferring the posterior exactly, we approximate the posterior using neural posterior estimation (NPE). NPE aims to select a variational approximation $q_{\phi}$ indexed by parameters $\phi\in\Phi$ that minimizes an expectation of the forward Kullback-Leibler (KL) divergence:
$\mathbb{E}_{(x,\pi,z)\sim p(x,\pi,z)}D_{\rm KL}(p(z\mid x,\pi)\,\|\,q_{\phi}(z\mid x,\pi))$
$=\mathbb{E}_{(x,\pi,z)\sim p(x,\pi,z)}\log q_{\phi}(z\mid x,\pi)+c,$
where $c$ is constant with respect to $\phi$.
We accomplish this by training a neural network to map $(x,\Pi)$ to a parameterized representation of $q(z\mid x,\Pi)$, which we describe further in Section 5.1. This network, whose architecture is described in Section 5.2, is trained through stochastic optimization using the simulated data described in Section 4. After the network has been trained, evaluating the posterior for a given data point $(x,\Pi)$ just involves a single forward pass.
@@ 5.1 The Variational Family
The probabilistic structure of our variational distribution, which factorizes over small spatial regions called “tiles,” matches that of Liu et al. (2023b). In particular, let $H^{\prime}=\lceil H/4\rceil$ and $W^{\prime}=\lceil W/4\rceil$. For $h=1,\ldots,H^{\prime}$ and $w=1,\ldots,W^{\prime}$, let image tile $T_{h,w}:=[h,h+4)\times[w,w+4)$ refer to a particular $\times4$-pixel region. Collectively, the image tiles are a partition of the region imaged by $x$, expressed in pixel coordinates. For a catalog $z$, let $z_{h,w}\subset z$ denote the cataloged light sources whose centroids are contained in $T_{h,w}$.
To approximate $p(z\mid x,\Pi)$, we use a family of variational distributions that factorize over tiles. In particular, we set
$q(z\mid x,\Pi):=\prod_{h=1}^{H^{\prime}}\prod_{w=1}^{W^{\prime}}q(z_{h,w}\mid x,\Pi),$
where the probabilistic structure of $q(z_{h,w}\mid x,\Pi)$, which we refer to as the per-tile variational distribution, is given in Table 2. For each tile, this distribution is governed by eight distribution parameters. For a tile with index $(h,w)$, these distributional parameters are outputted in spatial position $(h,w)$ by a fully convolutional inference network that takes $x$ as input. We describe its architecture next.
@@ 5.2 Neural Network Architecture
The input to our inference neural network is a sky-subtracted $80\times 80$-pixel image (i.e., $H=80$ and $W=80$) and a PSF (see Figure 3). We use a fully convolutional inference network to ensure strong translation equivariance, so sources produce consistent outputs regardless of where they appear. A straightforward way to incorporate the PSF would be to pad a rasterized version of it to match the input image resolution, then concatenate it to the image as an extra channel. This is the approach taken by Lanusse et al. (2021) in estimating galaxy morphologies from postage stamp images. However, estimating galaxy morphologies involves outputting a single prediction for the entire input image, whereas our network is responsible for predicting values for many sources appearing in different positions. For our problem, if we simply concatenate a rasterized PSF to the image, the network must integrate local information about individual source from the image channel while simultaneously needing long-range context from the PSF channel. Standard CNNs are optimized for localized processing and are not well-suited to integrate such long-range dependencies.
Instead, we provide the PSF to the inference network in its six-parameter form, which the SDSS pipeline fits to rasterized PSF images (see Section 2). Although the rasterized PSF generated the images, we now use the six‑parameter representation for inference. In Figure 3, we visualize this six-parameter PSF representation as produced by a generic PSF encoder, but at present, for simplicity, we take the PSF encoder to be the SDSS pipeline’s procedure for fitting six-parameter PSF models to rasterized PSFs. We discuss the possibility of using an auxiliary autoencoder as a more sophisticated method of obtaining this encoding in Section 7.
The two inputs—an $80\times{}80$-pixel input image and the length-6 PSF parameter vector—have different structures, so they cannot simply be stacked and provided as different channels to a CNN. Instead, we provide each of these six PSF parameters to the CNN with a dedicated feature map (channel) that is concatenated with the image along the channel axis. Each feature maps consists of the same value at each position and has the same height and width as the image. With this formulation, the PSF information can be accessed in the same way by a CNN at each pixel position.
6 Numerical Experiments
Through numerical experiments, we aim to assess the tradeoffs between training inference networks on narrowly focused versus diverse data distributions. Specifically, we compare networks trained on a single PSF and background—specialized to a particular observational setting—to those trained on a broad range of PSFs and backgrounds. Our main question is whether a generalist network, trained across heterogeneous conditions, can match the performance of a specialist network when evaluated on the specialist’s own domain. If so, this would obviate the need to train bespoke networks for each observed PSF–background combination.
We assess the performance of various inference network training regimes in terms of precision and recall in detection (Section 6.1.1), star/galaxy separation accuracy (Section 6.1.2), and flux measurement error (Section 6.1.3).